Don Bogdanov and the Russian bonds

Problem 1: Don Bogdanov trades international bonds at Donald & Low, a hedge fund based in NYC. Don is convinced that relations between Russia and Ukraine will improve over the near future.
He believes that, as a result, the Russian currency, the Ruble (RUB) will appreciate against the US dollar (USD), and that the price of Russian government bonds will also increase.

In the foreign exchange market, 1 USD currently trades at 34.20 RUB (bid) and 34.80 RUB (ask). In other words, the RUB/USD exchange rate = 34.20/34.80.
In the bond market, the 1-year Russian government bond currently trades at 80% of its par value (RUB 1,000,000). Don expects the market price of the bond to come back to its par value in a matter of weeks.

He decides to buy 100 of these bonds at their current price (80% of par) and hold them for 60 days.
In order to purchase the Russian bonds (denominated in RUB), Don must first purchase the right amount of RUB. To finance the purchase of these RUB in the foreign exchange market, he uses his firm
’s equity(denominated in USD, obviously) for 10% of the total amount and takes out a loan in USD for the remaining 90%. This USD loan has a 60-day maturity and carries a 3% per annum interest rate (interest paid at maturity, on day 60). The hedge fund internally charges an 8% per annum cost of equity for all trades using any amount of its equity. At the end of the 60-day holding period, he resells the bonds, now trading at 95% of par, and converts his RUB proceeds back into US Dollars, as the RUB/USD trades at 32.10/32.70 (bid/ask). Assume a 360-day year (12 months of 30 days each)

1) How many US dollars does Don need to borrow to initiate the trade?

2) What will be Don’s net profit (or loss) in USD on this trade, after all costs are taken into account?

3) Irrespective of your answer in Q.2, what would be the RUB/USD exchange rate at which Don would have

to resell his RUB at the 60-day maturity in order to break even on this deal, all costs included and all else remaining equal?
Answer #1

Given current exchange rate : 1 USD = 34.20 (bid) / 34.80 (ask) that is Don will have to convert USD into RUB at 34.20 (bid) . He purchases 100 bonds with face value RUB 1,000,000 at 80% par value. Hence his purchase price in RUB is (1000000*80%*100) = RUB 80,000,000.00 which in USD terms will be (80,000,000 / 34.20) = USD 2,339,181.29.

To fund this purchase, Don is using 10% of his firm's equity which is (2,339,181.29 * 10%) = USD 233,918.13 and balance will be funded through borrowing of USD 2,105,263.16

Now to calculate his net profit (or loss) we will first calculate his total cost in USD terms:

Cost of equity is 8%, hence (233918.13 * 8%) * 2/12 (since it is only for 60 days) = USD 3118.91

Cost of debt is 3%, hence (2105263.16*3%)*2/12 = USD 10,526.32

Now we will calculate his receivables from the sale of bonds - 100 bonds of RUB 1,000,000 face value sold at 95% par value which in RUB terms = 100 * 1000000 * 95% = RUB 95,000,000.00 At this time the exchange rate has become 1 USD = 32.10 (bid) / 32.70 (ask) . Don will be able to convert his RUB in USD at ask rate which is going to be (95000000/32.70) = USD 2,905,198.78

Hence his net profit = 2905198.78 - (2105263.16 + 10526.32 + 233918.13 + 3118.91) = USD 552,372.27

Now in case we want to calculate the breakeven exchange rate at the end of 60 days, then we have RUB receivable which is RUB 95 million and we have costs in USD which is USD 2,352,826.51 (sum of initial debt and equity portion and their respective costs thereof). The breakeven exhange rate will be (95000000/2352826.51) = 40.3770